OpenStudy (anonymous):

In an after school program, students must choose to either play on the playground or read a book silently. On one cloudy day, 70% of the students chose to play on the playground, and the rest chose read a book silently. 60% of the students who went to the playground are boys, and 20% who stayed inside to read a book are boys. If a student is randomly selected on this day, what is the probability that the student is a girl? A) 0.46 B) 0.48 C) 0.50 D) 0.52

OpenStudy (anonymous):

@ganeshie8 @Rushwr

OpenStudy (rushwr):

What did u get ?

OpenStudy (rushwr):

@CrazyCountryGirl

OpenStudy (rushwr):



OpenStudy (rushwr):

Girls who went to the playground are 40% of the student who went to the playground so probability of selecting a girl from the playground is \[\frac{ 70 }{ 100} * \frac{ 40 }{ 100 } = 0.28\] Girls who went read a book are 80% from the total of the students who reads books. So the probability to choose a girls from the reading section is \[\frac{ 30 }{ 100} * \frac{ 80 }{ 100 }\] that is 0.24 We are asked to find the total probability so we have to add those 2 up ! 0.24 + 0.28 =0.52

OpenStudy (badhi):

Same process of rushwr can be shown with following notation, \[P[G|Gr] = 0.4\\ P[G|B] =0.8\] \[P[B] = 0.3 \\P[Gr] = 0.7\] \[P[G] = P[G\cap Gr] + P[G \cap B]= P[G|Gr]\times P[Gr] + P[G|B]+P[B]\] G= Girls, Gr = in ground , B = book reading